# Upper and Lower Bounds for the Linear Ordering Principle

### v2

@@ Abstract
Korten and Pitassi (FOCS, 2024) defined a new [^1] Note that this notation had been used in the past [Sch83] for a very different class, which has been apparently forgotten after that. complexity class $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ as the polynomial-time Turing closure of the Linear Ordering Principle. They put it between $\mathbf{{MA}}$ (Merlin–Arthur protocols) and $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ (the second symmetric level of the polynomial hierarchy).
@@ 1.1.1 Classes Based on Symmetric Alternation
An important feature of these new results was that they were based on reducing finding a hard function to a total search problem. Namely, the works of Korten [Kor22] and Li [Li24] reduced the question to the so-called Range Avoidance problem: given a function $f\colon\{0,1\}^{n}\to\{0,1\}^{m}$ with $m>n$, represented by a Boolean circuit, find a point outside its image. In [CHR24, Li24], Range Avoidance has been reduced to symmetric alternation. Subsequently, Korten and Pitassi [KP24] reduced Range Avoidance to the Linear Ordering Principle: given an implicitly described ordering relation, either find the smallest element or report a breach of the linear order axioms. A polynomial-time Turing closure of this principle gave rise to a new class $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$: a version of $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ where the two provers provide points of a polynomial-time verifiable linear order on binary strings of a certain length (each point starting with the corresponding answer 0 or $1$), and the prover that provides the smaller element wins.
@@ 1.4.1 Approximate Counting, Set Size Estimation and $\mathbf{{SBP}}$
Computing the number of accepting paths of a given non-deterministic Turing machine is a fundamental problem captured by the “counting” class $\mathbf{\#\mathbf{{P}}}$. Yet, this class appears to be too powerful since, by Toda’s Theorem [Tod91], even a single query to it suffices to decide any language in the polynomial hierarchy, $\mathbf{{PH}}\subseteq\mathbf{{P}}^{\mathbf{\#\mathbf{{P}}[1]}}$! Given that, it is natural to explore approximations. To this end, one can consider the problem of Approximate Counting (ApproxCount for short) which refers to the task of approximating the number of accepting paths (within a constant factor). Equivalently, this problem can be framed as approximating the size of a set $S$ represented as the set of satisfying assignments of a Boolean circuit $C$. Previously, it was shown that this task could be carried out by a randomized algorithm using an $\mathbf{{NP}}$ oracle ($\mathbf{{FBPP}}^{\mathbf{{NP}}}$) [Sto85, JVV86] and by a deterministic algorithm using a $\mathbf{{prAM}}$ oracle ($\mathbf{{FP}}^{\mathbf{{prAM}}}$) [Sip83, GS86]. Shaltiel and Umans [SU06] show how to accomplish this task in $\mathbf{{FP}}^{\mathbf{{NP}}}$, yet under a derandomization assumption. We note that all of these algorithms can be implemented using parallel (i.e. non-adaptive) oracle queries. That is, in $\mathbf{{FBPP}}^{\mathbf{{NP}}}_{\parallel},\mathbf{{FP}}^{\mathbf{{prAM}}}_{\parallel}$ and $\mathbf{{FP}}^{\mathbf{{NP}}}_{\parallel}$, respectively. Approximate counting has also recently attracted considerable attention in the quantum literature [OS18, AKKT20, AR20, MAD25].
@@ 2.2 Problems Avoid and LOP
Korten [Kor22] proved that for a stretch of $n+1$ this problem is equivalent to a stretch of $O(n)$ (and, of course, vice versa) using $\mathbf{{P}}^{\mathbf{{NP}}}$ reductions.
The following definition is due to Korten and Pitassi [KP24]:
Definition 2.8 (LOP, Linear Ordering Principle).
@@ 3.1 Approximate Counting
  • For a non-empty subset $S\subseteq U$, $\left|{{\left\{(\upsilon,\alpha)\in U\times S\;\left|\;\upsilon<\alpha\right.\right\}}}\right|=\left|{S}\right|\cdot\mathrm{rank}(S).$
  • For any $\alpha\in U$, $\mathrm{rank}{{\left\{\upsilon\;\left|\;\upsilon\leq\alpha\right.\right\}}}=\mathrm{rank}(\alpha)/2$.
@@ 3.3 Finding the Minimum Using a $\mathbf{{prSBP}}$ Oracle
      /\$\$/˜If one of the sets is empty, we choose the other one
@@
      /\$\$/˜Choose the set with smaller approximate order
@@
Given Observation 3.8, we can assume w.l.o.g. that $E$ is a linear order circuit. We claim that the algorithm will output the minimal element after at most $2n$ iterations. Indeed, by Lemma 3.12, the order rank of the element $\alpha$ after $2n$ iterations satisfies $\mathrm{rank}(\alpha)\leq\frac{\mathrm{rank}(1^{n})}{\sqrt{2}^{2n}}\leq\frac{2^{n}-1}{2^{n}}<1$, thus the “While” cycle will be terminated before that. ∎
@@ 4.1 A Karp–Lipton–style Collapse to $\mathbf{{P}}^{\mathbf{{prOMA}}}$
Formally, a new Arthur $A^{\prime}$ expects $S_{n}$ as a proof, applies $S_{n}$ to $A_{x,r}$ itself and runs $A$ on the resulting witness. Therefore, we get exactly the definition of $\mathbf{{prOMA}}$ (Def. 2.13):
@@ 5.2 The Input-Oblivious Setting
Korten proves (Prop. 5.2) that PRG (Def. 5.1) reduces to Avoid, and Korten and Pitassi [KP24] compute Avoid in $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}=\mathbf{{P}}^{\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$. Since PRG has a unary input, one can observe that PRG can be computed using an input-oblivious oracle. This gives raise to tighter containments. In the non-input oblivious setting they are equalities, but input-oblivious classes lack some of the nice closureness properties and need a separate treatment.
@@
[…]
@@ 6 Discussion and Further Research
    Chakaravarthy and Roy [CR11] asked whether $\mathbf{{P}}^{\mathbf{{prMA}}}$ and $\mathbf{{P}}^{\mathbf{{prS_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$ are contained in $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. While we resolved the first question, the second one remains open. We note that, although in the input-oblivious world both inclusions hold ($\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, Corollary 4.4), the proof of the latter inclusion (Theorem 5) is essentially input-oblivious (one needs to give all the certificates for the oracle non-adaptively, and queries cannot be predicted because oracle answers cannot be predicted for promise problems).
@@
    We defined an input-oblivious version $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ of $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, however, contrary to $\mathbf{{NP}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, it is not obvious whether $\mathbf{{ONP}}\subseteq\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ or even $\mathbf{{ONP}}\subseteq\mathbf{{P}}^{\mathbf{{prOL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$.
@@
Acknowledgement
The authors are grateful to Yaroslav Alekseev for discussing and to Dmitry Itsykson for discussing and proofreading a preliminary version of this paper.

### v3

@@ Abstract
Korten and Pitassi (FOCS, 2024) defined a new [^1] Note that this notation had been used in the past [Sch83] for a very different class, which has been apparently forgotten after that. complexity class $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ as the polynomial-time Turing closure of the Linear Ordering Principle (a total function extending finding the minimum of an order [CK98] to the case where the order is not linear). They put it between $\mathbf{{MA}}$ (Merlin–Arthur protocols) and $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ (the second symmetric level of the polynomial hierarchy).
@@ 1.1.1 Classes Based on Symmetric Alternation
An important feature of these new results was that they were based on reducing finding a hard function to a total search problem. Namely, the works of Korten [Kor22] and Li [Li24] reduced the question to the so-called Range Avoidance problem: given a function $f\colon\{0,1\}^{n}\to\{0,1\}^{m}$ with $m>n$, represented by a Boolean circuit, find a point outside its image. This problem is known in the bounded arithmetic community as the dual Weak Pigeonhole Principle (dWPHP) and has notable implications in proof complexity (see e.g. [PWW88, Jeř04] and the survey [Kra25]). In [CHR24, Li24], Range Avoidance has been reduced to symmetric alternation. Subsequently, Korten and Pitassi [KP24] reduced Range Avoidance to the Linear Ordering Principle: given an implicitly described ordering relation, either find the smallest element or report a breach of the linear order axioms (for the case of a linear order this problem is known as MIN in the bounded arithmetic community [CK98]). A polynomial-time Turing closure of this principle gives rise to a new class $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$: a version of $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ where the two provers provide points of a polynomial-time verifiable linear order on binary strings of a certain length (each point starting with the corresponding answer 0 or $1$), and the prover that provides the smaller element wins.
@@ 1.4.1 Approximate Counting, Set Size Estimation and $\mathbf{{SBP}}$
Computing the number of accepting paths of a given non-deterministic Turing machine is a fundamental problem captured by the “counting” class $\mathbf{\#\mathbf{{P}}}$. Yet, this class appears to be too powerful since, by Toda’s Theorem [Tod91], even a single query to it suffices to decide any language in the polynomial hierarchy, $\mathbf{{PH}}\subseteq\mathbf{{P}}^{\mathbf{\#\mathbf{{P}}[1]}}$! Given that, it is natural to explore approximations. To this end, one can consider the problem of Approximate Counting (ApproxCount for short) which refers to the task of approximating the number of accepting paths (within a constant factor). Equivalently, this problem can be framed as approximating the size of a set $S$ represented as the set of satisfying assignments of a Boolean circuit $C$. Previously, it was shown that this task could be carried out by a randomized algorithm using an $\mathbf{{NP}}$ oracle ($\mathbf{{FBPP}}^{\mathbf{{NP}}}$) [Sto85, JVV86] and by a deterministic algorithm using a $\mathbf{{prAM}}$ oracle ($\mathbf{{FP}}^{\mathbf{{prAM}}}$) [Sip83, GS86]. Shaltiel and Umans [SU06] show how to accomplish this task in $\mathbf{{FP}}^{\mathbf{{NP}}}$, yet under a derandomization assumption. We note that all of these algorithms can be implemented using parallel (i.e. non-adaptive) oracle queries. That is, in $\mathbf{{FBPP}}^{\mathbf{{NP}}}_{\parallel},\mathbf{{FP}}^{\mathbf{{prAM}}}_{\parallel}$ and $\mathbf{{FP}}^{\mathbf{{NP}}}_{\parallel}$, respectively. In [Jeř07], Jeřábek studied approximate counting in the context of bounded arithmetic and reduced to it many problems in various complexity classes. Approximate counting has also recently attracted considerable attention in the quantum literature [OS18, AKKT20, AR20, MAD25].
@@ 2.2 Problems Avoid and LOP
Remark 2.8.
This problem, in a slightly different formulation, is known in the bounded arithmetic community as the dual Weak Pigeonhole Principle (dWPHP). See e.g. [Kra25] for more details.
Korten [Kor22] have shown that for a stretch of $n+1$ this problem is equivalent to a stretch of $O(n)$ (and, of course, vice versa) under $\mathbf{{P}}^{\mathbf{{NP}}}$ reductions.
The following definition is due to Korten and Pitassi [KP24]. It extends the search problem MIN introduced in [CK98], to the setting where the input relation is not necessarily a linear order.
Definition 2.9 (LOP, Linear Ordering Principle, [KP24]).
@@ 2.3 Complexity Classes
We also define $\mathbf{{SBP}}$ for the sake of self-completeness.
Definition 2.18 ([BGM06]).
A language $L$ is in $\mathbf{{SBP}}$ if there exist $\varepsilon>0$, $k,\ell\in\mathbb{N}$ and a polynomial-time computable predicate $B(x,r)$ such that
  • If $x\in L$ then $\implies\Pr_{r}[B(x,r)=1]\geq(1+\varepsilon)\cdot\frac{1}{2^{n^{k}}}$,
  • If $x\not\in L$ then $\implies\Pr_{r}[B(x,r)=1]\leq(1-\varepsilon)\cdot\frac{1}{2^{n^{k}}}$.
where $n=\left|{x}\right|$ and $r$ is uniformly distributed on $\{0,1\}^{n^{\ell}}$.
@@ 3.1 Approximate Counting
  • For a non-empty subset $S\subseteq U$: $\left|{{\left\{(\upsilon,\alpha)\in U\times S\;\left|\;\upsilon<\alpha\right.\right\}}}\right|=\left|{S}\right|\cdot\mathrm{rank}(S).$
  • For any $\alpha\in U$: $\mathrm{rank}{{\left\{\upsilon\in U\;\left|\;\upsilon\leq\alpha\right.\right\}}}=\mathrm{rank}(\alpha)/2$.
@@
Proof.
These observations follow from the definitions:
  • $\left|{{\left\{(\upsilon,\alpha)\in U\times S\;\left|\;\upsilon<\alpha\right.\right\}}}\right|=\sum\limits_{\alpha\in S}\left|{{\left\{(\upsilon,\alpha)\;\left|\;U\ni\upsilon<\alpha\right.\right\}}}\right|=\sum\limits_{\alpha\in S}\mathrm{rank}(\alpha)=\left|{S}\right|\cdot\mathrm{rank}(S).$
  • $\mathrm{rank}{{\left\{\upsilon\in U\;\left|\;\upsilon\leq\alpha\right.\right\}}}=\frac{1}{\mathrm{rank}(\alpha)+1}\cdot\sum\limits_{\upsilon\leq\alpha}{\mathrm{rank}(\upsilon)}=\frac{1}{\mathrm{rank}(\alpha)+1}\cdot\frac{\mathrm{rank}(\alpha)\left(\mathrm{rank}(\alpha)+1\right)}{2}=\frac{\mathrm{rank}(\alpha)}{2}$.
  • $\mathrm{rank}(S_{0}\cup S_{1})=\frac{1}{\left|{S_{0}}\right|+\left|{S_{1}}\right|}\cdot\left(\sum\limits_{x\in S_{0}}\mathrm{rank}(x)+\sum\limits_{y\in S_{1}}\mathrm{rank}(y)\right)=\frac{\left|{S_{0}}\right|\cdot\mathrm{rank}(S_{0})+\left|{S_{1}}\right|\cdot\mathrm{rank}(S_{1})}{\left|{S_{0}}\right|+\left|{S_{1}}\right|}$.
@@ 3.3 Finding the Minimum Using a $\mathbf{{prSBP}}$ Oracle
      /\$“!\$/˜If one of the sets is empty, we choose the other one
@@
      /\$“!\$/˜Choose the set with smaller approximate order
@@
Given Observation 3.8, we can assume w.l.o.g. that $E$ is a linear order circuit. We claim that the algorithm will output the minimal element after at most $2n$ iterations. Indeed, by Lemma 3.12, the order rank of the element $\alpha$ after $2n$ iterations satisfies $\mathrm{rank}(\alpha)\leq\frac{\mathrm{rank}(1^{n})}{\sqrt{2}^{2n}}\leq\frac{2^{n}-1}{2^{n}}<1$, thus the “While” cycle will terminate before that. ∎
@@ 4.1 A Karp–Lipton–style Collapse to $\mathbf{{P}}^{\mathbf{{prOMA}}}$
Formally, a new Arthur $A^{\prime}$ expects $S_{n}$ as a proof, applies $S_{n}$ to $A_{x,r}$ itself and runs $A$ on the resulting witness. Therefore, we get exactly the definition of $\mathbf{{prOMA}}$ (Def. 2.14):
@@ 4.3 Merging input-oblivious promise queries
Remark 4.6.
Unfortunately, the proof of Theorem 4.3 breaks down when one tries to extend it to $\mathbf{{P}}^{\mathbf{{prS_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$ since the queries to $\mathbf{{prS_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ are computed adaptively and the computation maybe be altered given the (unpredictable) answers to the queries that lie outside of promise.
However, almost the same proof shows that $\mathbf{{P}}^{\mathbf{{prS_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}_{||}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. Here all the queries can be computed from the input itself, thus instead of the input-oblivious witness $W_{n}$ one can give the non-input-oblivious witness specifically for these queries. The rest of the proof remains intact: $A^{\prime}$ will run $A$ on the witnesses intended for the particular query and proceed as before.
@@ 5.2 The Input-Oblivious Setting
Korten proves (Prop. 5.2) that PRG (Def. 5.1) reduces to Avoid, and Korten and Pitassi [KP24] compute Avoid in $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}=\mathbf{{P}}^{\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$. Since PRG has a unary input, one can observe that PRG can be computed using an input-oblivious oracle. This gives raise to tighter containments. Indeed, in the non-input oblivious setting this containments become equalities. However, input-oblivious classes lack some of the nice closure properties and therefore require a special treatment.
@@
The second inclusion in all the items follows from $\mathbf{{NP}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, $\mathbf{{prONP}}\subseteq\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ (syntactically), $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ (similarly to $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$), $\mathbf{{P}}^{\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}=\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [KP24]($\mathbf{{P}}^{\mathbf{{prL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$, because $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ is a syntactic class), $\mathbf{{P}}^{\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}=\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ ($\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$ (Theorem 4.3)). ∎
@@ 6 Discussion and Further Research
    Chakaravarthy and Roy [CR11] asked whether $\mathbf{{P}}^{\mathbf{{prMA}}}$ and $\mathbf{{P}}^{\mathbf{{prS_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$ are contained in $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. While we resolved the first question, the second one remains open We note that, although in the input-oblivious world both inclusions hold ($\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, Corollary 4.4), the proof of the latter inclusion (Theorem 5) is essentially input-oblivious (one needs to give all the certificates for the oracle non-adaptively, and queries cannot be predicted because oracle answers cannot be predicted for promise problems, thus in the non-input-oblivious setting it works for parallel queries only (Remark 4.6)).
@@
    We defined an input-oblivious version $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ of $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, it is not obvious whether $\mathbf{{ONP}}\subseteq\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ or even $\mathbf{{ONP}}\subseteq\mathbf{{P}}^{\mathbf{{prOL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$.
@@
Acknowledgment
The authors are grateful to Yaroslav Alekseev for discussions, to Dmitry Itsykson for discussing and proofreading a preliminary version of this paper, and to Jan Krajíček for providing multiple references and for useful discussions.